OpenStudy (madhatter):

6*-2=7-4y slove for y

jimthompson5910 (jim_thompson5910):

What does 6*-2 simplify to

OpenStudy (madhatter):

-12

jimthompson5910 (jim_thompson5910):

now you must move that 7 on the right side over, so how do you do this?

OpenStudy (madhatter):

-12+7=7+7-4y right?

jimthompson5910 (jim_thompson5910):

not adding, but subtracting

OpenStudy (madhatter):

oh ok so ....?????

jimthompson5910 (jim_thompson5910):

notice on the right side if you add 7 to both sides, you get 7+7 = 14 but we want 7 - 7 = 0 (so the 7 on the right side effectively goes away)

OpenStudy (madhatter):

OOOOOHHHHHHHH

jimthompson5910 (jim_thompson5910):

so you subtract 7 from both sides to get -12 = 7-4y -12-7 = 7-4y-7 -12-7 = 7-7-4y -19 = 0-4y -19 = -4y

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (madhatter):

mmmmmm -19/4=-4y/4

jimthompson5910 (jim_thompson5910):

good, that will give you -19/4 = -y flip the equation to get -y = -19/4 then change the signs of both sides to get this final answer y = 19/4

OpenStudy (madhatter):

4.75

jimthompson5910 (jim_thompson5910):

or you can leave it as a fraction

OpenStudy (madhatter):

whats the fraction?

jimthompson5910 (jim_thompson5910):

19/4

OpenStudy (madhatter):

oh duuuuhhhh lol

jimthompson5910 (jim_thompson5910):

lol that's ok

OpenStudy (madhatter):

thank so much ur the best

jimthompson5910 (jim_thompson5910):

sure thing